Kim owns a square piece of land with a side length of 'a' metres. He extends his property by purchasing adjacent land so that the length is increased by 𝟏𝟎 𝒎 and the width by 𝟏𝟐 𝒎. Write the simplified algebraic expression for
the area of Kim’s extended property. Provide a simplified response. You must solve by expanding. Thanks for the help. Appreciate it

Answer by ikleyn(52781)

.

The area of the extended property is

        (a+10)*(a+12) = a^2 + 10a + 12a + 120 = a^2 + 22a + 120 m^2     (simplified and completed).             ANSWER


--------------

It is  UNCLEAR  to me,  what difficulties did you met in solving this simple problem

and why you did not solve it on your own ?


Could you clarify   PLEASE ?